# Reinforced In-Context Black-Box Optimization

### v2

@@ 1 Introduction
Black-Box Optimization (BBO) [1, 4] refers to optimizing objective functions where neither analytic expressions nor derivatives of the objective are available. To solve BBO problems, we can only access the results of objective evaluation, which usually also incurs a high computational cost. Many fundamental problems in science and engineering involve optimization of expensive BBO functions, such as drug discovery [42, 55], material design [21, 23], robot control [10, 11], and optimal experimental design [24, 44], just to name a few.
@@
To date, a lot of BBO algorithms have been developed, among which the most prominent ones are Bayesian Optimization (BO) [20, 51] and Evolutionary Algorithms (EA) [5, 68]. Despite the advancements, these algorithms typically solve BBO problems from scratch and rely on expert-derived heuristics. Consequently, they are often hindered by slow convergence rates, and unable to leverage the inherent structures within the optimization problems [3, 6].
@@
Recently, there has been a growing interest in meta-learning a particular component of the algorithms with previously collected data [2, 18]. Learning the component not only alleviates the need for the laborious design process of the domain experts, but also specifies the component with domain data to facilitate subsequent optimization. For example, some components in BO are proposed to be learned from data, including the surrogate model [41, 46, 61, 64], acquisition function [29, 59], initialization strategy [17, 47], and search space [37, 45]; some core evolutionary operations in EA have also been considered, e.g., learning the selection and mutation rate adaptation in genetic algorithm [34] or the update rules for evolution strategy [35].
@@
There have also been some attempts to learn an entire algorithm in an End-to-End (E2E) fashion, which requires almost no expert knowledge at all and provides the most flexibility across a broad range of BBO problems. However, existing practices require additional knowledge regarding the objective function during the training stage, e.g., the gradient information (often impractical for BBO) [13, 56] or online sampling from the objective function (often very expensive) [39]Chen et al. [14] proposed the OptFormer method to imitate the behavior algorithms separately during training, requiring a user to manually specify which algorithm to execute during testing, which still poses a challenge. Thus, these methods are less ideal for practical scenarios where offline datasets are often available beforehand and a suitable algorithm for the given task has to be identified automatically without the involvement of domain experts.
@@
In this paper, we introduce Reinforced In-context BBO (RIBBO), which learns a reinforced BBO algorithm from offline datasets in an E2E fashion. RIBBO employs an expressive sequence model, i.e., causal transformer, to fit the optimization histories in the offline datasets generated by executing diverse behavior algorithms on multiple tasks. The sequence model is fed with previous query points and their function values, and trained to predict the distribution over the next query point. During testing, the sequence model itself serves as a BBO algorithm by generating the query points auto-regressively. Apart from this, RIBBO augments the optimization histories with regret-to-go (RTG) tokens, which are calculated by summing up the regrets over the future part of the histories, representing the future performance of an algorithm. A novel Hindsight Regret Relabelling (HRR) strategy is proposed to update the RTG tokens during testing. By integrating the RTG tokens into the modeling, RIBBO can automatically identify different algorithms, and generate sequences of query points that satisfy the specified regret. Such modeling enables RIBBO to circumvent the impact of inferior data and further reinforce its performance on top of the behavior algorithms.
@@
We perform experiments on BBOB synthetic functions, hyper-parameter optimization and robot control problems by using some representatives of heuristic search, EA, and BO as behavior algorithms to generate the offline datasets. The results show that RIBBO can automatically generate sequences of query points satisfying the user-desired regret across diverse problems, and achieve good performance universally. Note that the best behavior algorithm depends on the problem at hand, and RIBBO can perform even better on some problems. Compared to the most related method OptFormer [14], RIBBO also has clear advantage. In addition, we perform a series of experiments to analyze the influence of important components of RIBBO.
@@ 2.1 Black-Box Optimization
Let $f:\mathcal{X}\rightarrow\mathbb{R}$ be a black-box function, where $\mathcal{X}\subseteq\mathbb{R}^{d}$ is a $d$-dimensional search space. The goal of BBO is to find an optimal solution $\bm{x}^{*}\in\mathop{\arg\max}_{\bm{x}\in\mathcal{X}}f(\bm{x})$, with the only permission of querying the objective function value. Several classes of BBO algorithms have been proposed, e.g., BO [20, 51] and EA [5, 68]. The basic framework of BO contains two critical components: a surrogate model, typically formalized as Gaussian Process (GP) [49], and an acquisition function [63], which are used to model $f$ and decide the next query point, respectively. EA is a class of heuristic optimization algorithms inspired by natural evolution. It maintains a population of solutions and iterates through mutation, crossover, and selection operations to find better solutions.
@@ 2.2 Meta-Learning in Black-Box Optimization
Hand-crafted BBO algorithms usually require an expert to analyze the algorithms’ behavior on a wide range of problems, which is tedious and time-consuming. One solution is meta-learning [28, 58], which aims to exploit knowledge to improve the performance of learning algorithms given data from a set of tasks. By parameterizing a component of BBO algorithms or even an entire algorithm that is traditionally manually designed, we can learn from historical data to incorporate the knowledge into the optimization, which may bring speedup.
@@
Meta-learning particular components has been studied with different BBO algorithms. Meta-learning in BO can be divided into four main categories according to “what to transfer” [6], including the design of the surrogate model, acquisition function, initialization strategy, and search space. For surrogate model design, Wang et al. [61] and Wistuba and Grabocka [64] parameterized the mean or kernel function of the GP model with Multi-Layer Perceptron (MLP), while Perrone et al. [46] and Müller et al. [41] substituted GP with Bayesian linear regression or neural process [22, 40]. For acquisition function design, MetaBO [59] uses Reinforcement Learning (RL) to meta-train an acquisition function on a set of related tasks, and FSAF [29] employs a Bayesian variant of deep Q-network as a surrogate differentiable acquisition function trained by model-agnostic meta-learning [19]. The remaining two categories focus on exploiting the previous good solutions to warm start the optimization [17, 47] or shrink the search space [37, 45]. Meta-learning in EA usually focuses on specific evolutionary operations. For example, Lang et al. substituted core genetic operators, i.e., selection and mutation rate adaptation, with dot-product attention modules [34], and meta-learned a self-attention-based architecture to discover effective and order-invariant update rules [35].
@@
Meta-learning entire algorithms has also been explored to obtain more flexible models. Early works [13, 56] use Recurrent Neural Network (RNN) to meta-learn a BBO algorithm by optimizing the summed objective functions of some iterations. RNN uses its memory state to store information about history and outputs the next query point. This work assumes access to gradient information during the training phase, which is, however, usually impractical in BBO. OptFormer [14] uses a text-based transformer framework to learn an algorithm, providing a universal E2E interface for BBO problems. It can simultaneously imitate at least $7$ different BBO algorithms on a broad range of problems, which, however, requires a user to specify an algorithm manually when applying it to solve problems. Thus, how to select suitable algorithms automatically is still a challenge. Neural Acquisition Processes (NAP) [39] uses transformer to meta-learn the surrogate model and acquisition function of BO jointly. Due to the lack of labeled acquisition data, NAP uses an online RL algorithm with a supervised auxiliary loss for training, which requires online sampling from the expensive objective function and lacks efficiency. Black-box Optimization NETworks (BONET) [33] employ a transformer model to fit regret-augmented trajectories in an offline BBO scenario, where the training and testing data are from the same objective function, and a prefix sequence is required to warm up the optimization before testing. Compared to the above state-of-the-art E2E methods, we consider the meta-BBO setting, where the training datasets consist of diverse algorithms across different functions. Our approach offers the advantage of automatically identifying and deploying the best-performing algorithm without requiring the user to pre-specify which algorithm to use or to provide a prefix sequence during the testing phase, and utilizes a supervised learning loss for training on a fixed offline dataset without the need for further interaction with the objective function.
@@ 2.3 Decision Transformer
Transformer has emerged as a powerful architecture for sequence modeling tasks [32, 62, 65]. A basic building block behind transformer is the self-attention mechanism [57], which captures the correlation between tokens of any pair of timesteps. As the scale of data and model increases, transformer has demonstrated the in-context learning ability [9], which refers to the capability of the model to infer the tasks at hand based on the input contexts. Decision Transformer (DT) [12] abstracts RL as a sequence modeling problem, and introduces return-to-go tokens, representing the cumulative rewards over future interactions. Conditioning on return-to-go tokens enables DT to correlate the trajectories with their corresponding returns and generate future actions to achieve a user-specified return. Inspired by DT, we will treat BBO tasks as a sequence modeling problem naturally, use a causal transformer for modeling, and train it by conditioning on future regrets. Such design is expected to enable the learned model to distinguish algorithms with different performance and achieve good performance with a user-specified low regret.
@@ 3 Method
This section presents Reinforced In-context Black-Box Optimization (RIBBO), which learns an enhanced BBO algorithm in an E2E fashion, as illustrated in Figure 1. We follow the task-distribution assumption, which is commonly adopted in meta-learning settings [19, 28, 67]. Our goal is to learn a generalizable model $\mathcal{M}$ capable of solving a wide range of BBO tasks, each associated with a BBO objective function $f$ sampled from the task distribution $P(\mathcal{F})$, where $\mathcal{F}$ denotes the function space.
@@ 3.1 Method Outline
Given the current history $\bm{h}_{t-1}$ at iteration $t$, a BBO algorithm selects the next query point $\bm{x}_{t}$, observes the function value $y_{t}=f_{i}(\bm{x}_{t})$, and updates the history $\bm{h}_{t}=\bm{h}_{t-1}\cup\{(\bm{x}_{t},y_{t})\}$. Similar to the previous work [13], we take this framework as a starting point and treat the learning of a universal BBO algorithm as learning a model $\mathcal{M}_{\bm{\theta}}$, which takes the preceding history $\bm{h}_{t-1}$ as input and outputs a distribution of the next query point $\bm{x}_{t}$. The optimization histories in offline datasets provide a natural supervision for learning.
Given a set of histories $\{\bm{h}_{T}\}$ generated by a single behavior algorithm $\mathcal{A}$ on a single task $f$, we employ a causal transformer model $\mathcal{M}_{\bm{\theta}}$, and expect it to imitate $\mathcal{A}$ and produce similar optimization history on $f$. In practice, we usually have datasets containing histories from multiple behavior algorithms $\{\mathcal{A}_{j}\}_{j=1}^{K}$ on multiple tasks $\{f_{i}\}_{i=1}^{N}$. To fit $\mathcal{M}_{\bm{\theta}}$, we use the negative log-likelihood loss
@@
Nevertheless, naively imitating the offline datasets hinders the model since some inferior behavior algorithms may severely degenerate the model’s performance. Inspired by DT [12], we propose to augment the optimization history with Regret-To-Go (RTG) tokens $R_{t}$, defined as the sum of instantaneous regrets over the future history:
@@
The resulting method RIBBO has implicitly utilized the in-context learning capacity of transformer to guide the optimization with previous histories and the desired future regret as context. The in-context learning capacity of inferring the tasks at hand based on the input contexts has been observed as the scale of data and model increases [31]. It has been explored to infer general functional relationships as supervised learning or RL algorithms. For example, the model is expected to predict accurately on the query input $\bm{x}_{t}$ by feeding the training dataset $\{(\bm{x}_{i},y_{i})\}_{i=1}^{t-1}$ as the context [25, 27, 38]Laskin et al. [36] learned RL algorithms using causal transformers. Here, we use it for BBO.
@@ 3.2 Practical Implementation
Model Architecture. For the formalization of the model $\mathcal{M}_{\bm{\theta}}$, we adopt the commonly used GPT architecture [48], which comprises a stack of causal attention blocks, each containing an attention component and a feed-forward network. We aggregate each triplet $(\bm{x}_{i},y_{i},R_{i})$ using a two-layer MLP network. The output of $\mathcal{M}_{\bm{\theta}}$ is a diagonal Gaussian distribution of the next query point. Note that previous works that adopt the sequence model as a surrogate model [40, 41, 43] typically remove the positional encoding because the surrogate model should be invariant to the history order. On the contrary, our implementation preserves the positional encoding, naturally following the behavior of certain algorithms (e.g., BO or EA) and making it easier to learn from algorithms. Additionally, the positional encoding can keep the monotonically decreasing order of RTG tokens. More architecture details can be found in Appendix A.
Model Training. RTG tokens are computed as outlined in Eq. (3) for the offline datasets before training. As the calculation of regret requires the optimum value of task $i$, we use the best-observed value $y_{\rm max}^{i}$ as a proxy for the optimum. Let $\{\widehat{\mathcal{D}}_{i,j}\}_{i\in[N],j\in[K]}$ denote the RTG augmented datasets with $N$ tasks and $K$ algorithms. We sample a minibatch of consecutive subsequences of length $\tau<T$ uniformly from the augmented datasets and train $\mathcal{M}_{\bm{\theta}}$ to minimize the RIBBO loss in Eq. (4).
@@
Model Inference. The model $\mathcal{M}_{\bm{\theta}}$ generates the query points $\bm{x}_{t}$ auto-regressively during the inference process, which involves iteratively selecting a new query point $\bm{x}_{t}$ based on the current augmented history $\hat{\bm{h}}_{t-1}$, evaluating the query as $y_{t}=f(\bm{x}_{t})$, and updating the history $\hat{\bm{h}}_{t}=\hat{\bm{h}}_{t-1}\cup\{(\bm{x}_{t},y_{t},R_{t})\}$. A critical aspect of this process is how to specify the value of RTG (i.e., $R_{t}$) at every iteration $t$. Inspired by DT, a naive approach is to specify a desired performance as the initial RTG $R_{0}$, and decrease it as $R_{t}=R_{t-1}-(y^{*}-y_{t})$. However, this strategy has the risk of producing out-of-distribution RTGs, since the values can fall below 00 due to an improperly selected $R_{0}$.
Given the fact that RTGs are lower bounded by 00 and a 00 value implies a good BBO algorithm with low regret, we propose to set the immediate RTG as 0, and use a strategy called Hindsight Regret Relabelling (HRR) to update previous RTGs based on the current sampling. The inference procedure with HRR is detailed in Algorithm 1. In line 1, the history $\hat{\bm{h}}_{0}$ is initialized with padding placeholders $\bm{x}_{0},y_{0}$ and RTG $R_{0}=0$. At iteration $t$ (i.e., lines 3–7), the model $\mathcal{M}_{\bm{\theta}}$ is fed with the augmented history $\hat{\bm{h}}_{t-1}$ to generate the next query point $\bm{x}_{t}$ in line 3, which is evaluated to obtain $y_{t}$ in line 4. Then, the immediate RTG $R_{t}$ is set to 00, and we employ HRR to update RTG tokens in $\hat{\bm{h}}_{t-1}$, i.e., calculate the instantaneous regret $r=y^{*}-y_{t}$ (line 5) and add $r$ to every RTG token in $\hat{\bm{h}}_{t-1}$ (line 6):
@@
Note that this relabelling process guarantees that $\forall 0\leq i<t$, the RTG token $R_{i}=\sum_{t^{\prime}=i+1}^{t}(y^{*}-y_{t^{\prime}})$, which can also be written as $\sum_{t^{\prime}=i+1}^{T}(y^{*}-y_{t^{\prime}})$, consistent with the definition in Eq. (3), because the immediate RTG $R_{t}=\sum_{t^{\prime}=t+1}^{T}(y^{*}-y_{t^{\prime}})$ is set to 0. In line 7, the history $\hat{\bm{h}}_{t}$ is updated by expanding $\hat{\bm{h}}_{t-1}$ with $\{(\bm{x}_{t},y_{t},0)\}$, i.e., the current sampling and its immediate RTG $R_{t}=0$. This process is repeated until reaching the budget $T$. Thus, we can find that HRR not only exploits the full potential of $\mathcal{M}_{\bm{\theta}}$ through using 00 as the immediate RTG and thereby demands the model to generate the most advantageous decisions, but also preserves the calculation of RTG tokens following the same way as the training data, i.e., representing the cumulative regret over future optimization history.

### v3

@@ 1 Introduction
Black-Box Optimization (BBO) Alarie et al. (2021); Audet & Hare (2017) refers to optimizing objective functions where neither analytic expressions nor derivatives of the objective are available. To solve BBO problems, we can only access the results of objective evaluation, which usually also incurs a high computational cost. Many fundamental problems in science and engineering involve optimization of expensive BBO functions, such as drug discovery Negoescu et al. (2011); Terayama et al. (2021), material design Frazier & Wang (2016); Gómez-Bombarelli et al. (2018), robot control Calandra et al. (2016); Chatzilygeroudis et al. (2019), and optimal experimental design Greenhill et al. (2020); Nguyen et al. (2023), just to name a few.
@@
To date, a lot of BBO algorithms have been developed, among which the most prominent ones are Bayesian Optimization (BO) Frazier (2018); Shahriari et al. (2016) and Evolutionary Algorithms (EA) Back (1996); Zhou et al. (2019). Despite the advancements, these algorithms typically solve BBO problems from scratch and rely on expert-derived heuristics. Consequently, they are often hindered by slow convergence rates, and unable to leverage the inherent structures within the optimization problems Astudillo & Frazier (2021); Bai et al. (2023).
@@
Recently, there has been a growing interest in meta-learning a particular component of the algorithms with previously collected data Arango et al. (2021); Feurer et al. (2021). Learning the component not only alleviates the need for the laborious design process of the domain experts, but also specifies the component with domain data to facilitate subsequent optimization. For example, some components in BO are proposed to be learned from data, including the surrogate model Müller et al. (2023); Perrone et al. (2018); Wang et al. (2021); Wistuba & Grabocka (2021), acquisition function Hsieh et al. (2021); Volpp et al. (2020), initialization strategy Feurer et al. (2015); Poloczek et al. (2016), and search space Li et al. (2022); Perrone & Shen (2019); some core evolutionary operations in EA have also been considered, e.g., learning the selection and mutation rate adaptation in genetic algorithm Lange et al. (2023a) or the update rules for evolution strategy Lange et al. (2023b); the configuration of the algorithm can also be learned and dynamically adjusted throughout the optimization process Biedenkapp et al. (2020); Adriaensen et al. (2022).
@@
There have also been some attempts to learn an entire algorithm in an End-to-End (E2E) fashion, which requires almost no expert knowledge at all and provides the most flexibility across a broad range of BBO problems. However, existing practices require additional knowledge regarding the objective function during the training stage, e.g., the gradient information (often impractical for BBO) Chen et al. (2017); TV et al. (2019) or online sampling from the objective function (often very expensive) Maraval et al. (2023)Chen et al. (2022) proposed the OptFormer method to imitate the behavior algorithms separately during training, presenting a challenge for the user to manually specify which algorithm to execute during testing. Thus, these methods are less ideal for practical scenarios where offline datasets are often available beforehand and a suitable algorithm for the given task has to be identified automatically without the involvement of domain experts.
@@
In this paper, we introduce Reinforced In-context BBO (RIBBO), which learns a reinforced BBO algorithm from offline datasets in an E2E fashion. RIBBO employs an expressive sequence model, i.e., causal transformer, to fit the optimization histories in the offline datasets generated by executing diverse behavior algorithms on multiple tasks. The sequence model is fed with previous query points and their function values, and trained to predict the distribution over the next query point. During testing, the sequence model itself serves as a BBO algorithm by generating the next query points auto-regressively. Apart from this, RIBBO augments the optimization histories with regret-to-go (RTG) tokens, which are calculated by summing up the regrets over the future part of the histories, representing the future performance of an algorithm. A novel Hindsight Regret Relabelling (HRR) strategy is proposed to update the RTG tokens during testing. By integrating the RTG tokens into the modeling, RIBBO can automatically identify different algorithms, and generate sequences of query points that satisfy the specified regret. Such modeling enables RIBBO to circumvent the impact of inferior data and further reinforce its performance on top of the behavior algorithms.
@@
We perform experiments on BBOB synthetic functions, hyper-parameter optimization and robot control problems by using some representatives of heuristic search, EA, and BO as behavior algorithms to generate the offline datasets. The results show that RIBBO can automatically generate sequences of query points satisfying the user-desired regret across diverse problems, and achieve good performance universally. Note that the best behavior algorithm depends on the problem at hand, and RIBBO can perform even better on some problems. Compared to the most related method OptFormer Chen et al. (2022), RIBBO also has clear advantage. In addition, we perform a series of experiments to analyze the influence of important components of RIBBO.
@@ 2.1 Black-Box Optimization
Let $f:\mathcal{X}\rightarrow\mathbb{R}$ be a black-box function, where $\mathcal{X}\subseteq\mathbb{R}^{d}$ is a $d$-dimensional search space. The goal of BBO is to find an optimal solution $\bm{x}^{*}\in\mathop{\arg\max}_{\bm{x}\in\mathcal{X}}f(\bm{x})$, with the only permission of querying the objective function value. Several classes of BBO algorithms have been proposed, e.g., BO Frazier (2018); Shahriari et al. (2016) and EA Back (1996); Zhou et al. (2019). The basic framework of BO contains two critical components: a surrogate model, typically formalized as Gaussian Process (GP) Rasmussen & Williams (2006), and an acquisition function Wilson et al. (2018), which are used to model $f$ and decide the next query point, respectively. EA is a class of heuristic optimization algorithms inspired by natural evolution. It maintains a population of solutions and iterates through mutation, crossover, and selection operations to find better solutions.
@@ 2.2 Meta-Learning in Black-Box Optimization
Hand-crafted BBO algorithms usually require an expert to analyze the algorithms’ behavior across a wide range of problems, a process that is both tedious and time-consuming. One solution is meta-learning Vilalta & Drissi (2002); Hospedales et al. (2021), which aims to exploit knowledge to improve the performance of learning algorithms given data from a collection of tasks. By parameterizing a component of BBO algorithms or even an entire BBO algorithm that is traditionally manually designed, we can utilize historical data to incorporate the domain knowledge into the optimization, which may bring speedup.
@@
Meta-learning particular components has been studied with different BBO algorithms. Meta-learning in BO can be divided into four main categories according to “what to transfer” Bai et al. (2023), including the design of the surrogate model, acquisition function, initialization strategy, and search space. For surrogate model design, Wang et al. (2021) and Wistuba & Grabocka (2021) parameterized the mean or kernel function of the GP model with Multi-Layer Perceptron (MLP), while Perrone et al. (2018) and Müller et al. (2023) substituted GP with Bayesian linear regression or neural process Garnelo et al. (2018); Müller et al. (2022). For acquisition function design, MetaBO Volpp et al. (2020) uses Reinforcement Learning (RL) to meta-train an acquisition function on a set of related tasks, and FSAF Hsieh et al. (2021) employs a Bayesian variant of deep Q-network as a surrogate differentiable acquisition function trained by model-agnostic meta-learning Finn et al. (2017). The remaining two categories focus on exploiting the previous good solutions to warm start the optimization Feurer et al. (2015); Poloczek et al. (2016) or shrink the search space Perrone & Shen (2019); Li et al. (2022). Meta-learning in EA usually focuses on learning specific evolutionary operations. For example, Lang et al. substituted core genetic operators, i.e., selection and mutation rate adaptation, with dot-product attention modules Lange et al. (2023a), and meta-learned a self-attention-based architecture to discover effective and order-invariant update rules Lange et al. (2023b). ALDes Zhao et al. (2024) introduces an auto-regressive learning-based approach to sequentially generate components of meta-heuristic algorithms. Beyond that, Dynamic Algorithm Configuration (DAC) Biedenkapp et al. (2020); Adriaensen et al. (2022) concentrates on learning the configurations of algorithms, employing RL to dynamically adjust the configurations during the optimization process.
@@
Meta-learning entire algorithms has also been explored to obtain more flexible models. Early works Chen et al. (2017); TV et al. (2019) use Recurrent Neural Network (RNN) to meta-learn a BBO algorithm by optimizing the summed objective functions of some iterations. RNN uses its memory state to store information about history and outputs the next query point. This work assumes access to gradient information during the training phase, which is, however, usually impractical in BBO problems. OptFormer Chen et al. (2022) uses a text-based transformer framework to learn an algorithm, providing a universal E2E interface for BBO problems. It is trained to imitate different BBO algorithms across a broad range of problems, which, however, presents a challenge for the user to manually specify an algorithm for inference. Neural Acquisition Processes (NAP) Maraval et al. (2023) uses transformer to meta-learn the surrogate model and acquisition function of BO jointly. Due to the lack of labeled acquisition data, NAP uses an online RL algorithm with a supervised auxiliary loss for training, which requires online sampling from the expensive objective function and lacks efficiency. Black-box Optimization NETworks (BONET) Krishnamoorthy et al. (2023) employ a transformer model to fit regret-augmented trajectories in an offline BBO scenario, where the training and testing data are from the same objective function, and a prefix sequence is required to warm up the optimization before testing. Compared to the above state-of-the-art E2E methods, we consider the meta-BBO setting, where the training datasets consist of diverse algorithms across different functions. Our approach offers the advantage of automatically identifying with RTG tokens and deploying the best-performing algorithm without requiring the user to pre-specify which algorithm to use or to provide a prefix sequence during the testing phase. It utilizes a supervised learning loss for training on a fixed offline dataset without the need for further interaction with the objective function.
@@ 2.3 Decision Transformer
Transformer has emerged as a powerful architecture for sequence modeling tasks Khan et al. (2022); Wen et al. (2022); Wolf et al. (2020). A basic building block behind transformer is the self-attention mechanism Vaswani et al. (2017), which captures the correlation between tokens of any pair of timesteps. As the scale of data and model increases, transformer has demonstrated the in-context learning ability Brown et al. (2020), which refers to the capability of the model to infer the tasks at hand based on the input contexts. Decision Transformer (DT) Chen et al. (2021) abstracts RL as a sequence modeling problem, and introduces return-to-go tokens, representing the cumulative rewards over future interactions. Conditioning on return-to-go tokens enables DT to correlate the trajectories with their corresponding returns and generate future actions to achieve a user-specified return. Inspired by DT, we will treat BBO tasks as a sequence modeling problem naturally, use a causal transformer for modeling, and train it by conditioning on future regrets. Such design is expected to enable the learned model to distinguish algorithms with different performance and achieve good performance with a user-specified low regret.
@@ 3 Method
This section presents Reinforced In-context Black-Box Optimization (RIBBO), which learns an enhanced BBO algorithm in an E2E fashion, as illustrated in Figure 1. We follow the task-distribution assumption, which is commonly adopted in meta-learning settings Finn et al. (2017); Hospedales et al. (2021); Zhou et al. (2023). Our goal is to learn a generalizable model $\mathcal{M}$ capable of solving a wide range of BBO tasks, each associated with a BBO objective function $f$ sampled from the task distribution $P(\mathcal{F})$, where $\mathcal{F}$ denotes the function space.
@@ 3.1 Method Outline
Given the current history $\bm{h}_{t-1}$ at iteration $t$, a BBO algorithm selects the next query point $\bm{x}_{t}$, observes the function value $y_{t}=f_{i}(\bm{x}_{t})$, and updates the history $\bm{h}_{t}=\bm{h}_{t-1}\cup\{(\bm{x}_{t},y_{t})\}$. Similar to the previous work Chen et al. (2017), we take this framework as a starting point and treat the learning of a universal BBO algorithm as learning a model $\mathcal{M}_{\bm{\theta}}$, which takes the preceding history $\bm{h}_{t-1}$ as input and outputs a distribution of the next query point $\bm{x}_{t}$. The optimization histories in offline datasets provide a natural supervision for the learning process.
Suppose we have a set of histories $\{\bm{h}_{T}\}$, generated by a single behavior algorithm $\mathcal{A}$ on a single task $f$. By employing a causal transformer model $\mathcal{M}_{\bm{\theta}}$, we expect $\mathcal{M}_{\bm{\theta}}$ to imitate $\mathcal{A}$ and produce similar optimization history on $f$. In practice, we usually have datasets containing histories from multiple behavior algorithms $\{\mathcal{A}_{j}\}_{j=1}^{K}$ on multiple tasks $\{f_{i}\}_{i=1}^{N}$. To fit $\mathcal{M}_{\bm{\theta}}$, we use the negative log-likelihood loss
@@
Nevertheless, naively imitating the offline datasets hinders the model since some inferior behavior algorithms may severely degenerate the model’s performance. Inspired by DT Chen et al. (2021), we propose to augment the optimization history with Regret-To-Go (RTG) tokens $R_{t}$, defined as the sum of instantaneous regrets over the future history:
@@
The resulting method RIBBO has implicitly utilized the in-context learning capacity of transformer to guide the optimization with previous histories and the desired future regret as context. The in-context learning capacity of inferring the tasks at hand based on the input contexts has been observed as the scale of data and model increases Kaplan et al. (2020). It has been explored to infer general functional relationships as supervised learning or RL algorithms. For example, the model is expected to predict accurately on the query input $\bm{x}_{t}$ by feeding the training dataset $\{(\bm{x}_{i},y_{i})\}_{i=1}^{t-1}$ as the context Guo et al. (2023); Hollmann et al. (2023); Li et al. (2023)Laskin et al. (2023) learned RL algorithms using causal transformers. Here, we use it for BBO.
@@ 3.2 Practical Implementation
Model Architecture. For the formalization of the model $\mathcal{M}_{\bm{\theta}}$, we adopt the commonly used GPT architecture Radford et al. (2018), which comprises a stack of causal attention blocks. Each block is composed of an attention mechanism and a feed-forward network. We aggregate each triplet $(\bm{x}_{i},y_{i},R_{i})$ using a two-layer MLP network. The output of $\mathcal{M}_{\bm{\theta}}$ is a diagonal Gaussian distribution of the next query point. Note that previous works that adopt the sequence model as surrogate models Müller et al. (2022; 2023); Nguyen & Grover (2022) typically remove the positional encoding because the surrogate model should be invariant to the history order. On the contrary, our implementation preserves the positional encoding, naturally following the behavior of certain algorithms (e.g., BO or EA) and making it easier to learn from algorithms. Additionally, the positional encoding can help maintain the monotonically decreasing order of RTG tokens. More details about the architecture can be found in Appendix A.
Model Training. RTG tokens are calculated as outlined in Eq. (3) for the offline datasets before training. Since the calculation of regret requires the optimum value of task $i$, we use the best-observed value $y_{\rm max}^{i}$ as a proxy for the optimum. Let $\{\widehat{\mathcal{D}}_{i,j}\}_{i\in[N],j\in[K]}$ denote the RTG augmented datasets with $N$ tasks and $K$ algorithms. During training, we sample a minibatch of consecutive subsequences of length $\tau<T$ uniformly from the augmented datasets. The training objective is to minimize the RIBBO loss in Eq. (4).
@@
Model Inference. The model $\mathcal{M}_{\bm{\theta}}$ generates the query points $\bm{x}_{t}$ auto-regressively during inference, which involves iteratively selecting a new query point $\bm{x}_{t}$ based on the current augmented history $\hat{\bm{h}}_{t-1}$, evaluating the query as $y_{t}=f(\bm{x}_{t})$, and updating the history by $\hat{\bm{h}}_{t}=\hat{\bm{h}}_{t-1}\cup\{(\bm{x}_{t},y_{t},R_{t})\}$. A critical aspect of this process is how to specify the value of RTG (i.e., $R_{t}$) at every iteration $t$. Inspired by DT, a naive approach is to specify a desired performance as the initial RTG $R_{0}$, and decrease it as $R_{t}=R_{t-1}-(y^{*}-y_{t})$. However, this strategy has the risk of producing out-of-distribution RTGs, since the values can fall below 00 due to an improperly selected $R_{0}$.
Given the fact that RTGs are lower bounded by 00 and a value of 00 implies a good BBO algorithm with low regret, we propose to set the immediate RTG as 00. Furthermore, we introduce a strategy called Hindsight Regret Relabelling (HRR) to update previous RTGs based on the current sample evaluations. The inference procedure with HRR is detailed in Algorithm 1. In line 1, the history $\hat{\bm{h}}_{0}$ is initialized with padding placeholders $\bm{x}_{0},y_{0}$ and RTG $R_{0}=0$. At iteration $t$ (i.e., lines 3–7), the model $\mathcal{M}_{\bm{\theta}}$ is fed with the augmented history $\hat{\bm{h}}_{t-1}$ to generate the next query point $\bm{x}_{t}$ in line 3, followed by the evaluation procedure to obtain $y_{t}$ in line 4. Then, the immediate RTG $R_{t}$ is set to 00, and we employ HRR to update previous RTG tokens in $\hat{\bm{h}}_{t-1}$, i.e., calculate the instantaneous regret $r=y^{*}-y_{t}$ (line 5) and add $r$ to every RTG token within $\hat{\bm{h}}_{t-1}$ (line 6):
@@
Note that this relabelling process guarantees that $\forall 0\leq i<t$, the RTG token $R_{i}=\sum_{t^{\prime}=i+1}^{t}(y^{*}-y_{t^{\prime}})$, which can also be written as $\sum_{t^{\prime}=i+1}^{T}(y^{*}-y_{t^{\prime}})$, consistent with the definition in Eq. (3), because the immediate RTG $R_{t}=\sum_{t^{\prime}=t+1}^{T}(y^{*}-y_{t^{\prime}})$ is set to 0. In line 7, the history $\hat{\bm{h}}_{t}$ is updated by expanding $\hat{\bm{h}}_{t-1}$ with $\{(\bm{x}_{t},y_{t},0)\}$, i.e., the current sampling and its immediate RTG $R_{t}=0$. The above process is repeated until reaching the budget $T$. Thus, we can find that HRR not only exploits the full potential of $\mathcal{M}_{\bm{\theta}}$ through using 00 as the immediate RTG and thereby demands the model to generate the most advantageous decisions, but also preserves the calculation of RTG tokens following the same way as the training data, i.e., representing the cumulative regret over future optimization history.